# Finite groups and arc-transitive maps of square-free Euler characteristic

### v2

@@ 4. $p$-Groups
Let $(x,y,z)$ be a regular triple for $G$. If $\overline{x},\overline{y},\overline{z}\in\overline{G}$ are all involutions, and $\overline{x}\neq\overline{z}$, then they form a regular triple for $\overline{G}$, since $\overline{G}=\langle\overline{x},\overline{y},\overline{z}\rangle$ and $\overline{x}\,\overline{z}=\overline{z}\,\overline{x}$. Otherwise, either some of $\overline{x},\overline{y},\overline{z}$ equals the identity-element in $\overline{G}$, or $\overline{x}=\overline{z}$. In both cases, $\overline{G}$ is generated by one or two involutions and then it is ${\sf Z}_{2}$ or dihedral.
@@
Let $G={\rm Q}_{8}{:}{\rm S}_{3}\cong{\rm GL}(2,3)$. Then neither $G$ nor ${\sf Z}_{4}\circ G$ has a regular triple.
@@
Suppose that $G$ has a regular triple $(x,y,z)$. Since $xz=zx$, we have $\langle x,z\rangle\cong{\sf Z}_{2}^{2}$, and so $\{x,z,xz\}\cap{\rm Q}_{8}\not$$\emptyset$. That is, one of $x,z,xz$ has to be the central involution 1-1 of $G$. It follows that $\overline{G}=G/\langle-1\rangle=\langle\overline{x},\overline{y}\rangle$ or $\langle\overline{y},\overline{z}\rangle$ is a dihedral group, which is impossible as $\overline{G}={\sf Z}_{2}^{2}{:}{\rm S}_{3}$.
Suppose that $K={\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm S}_{3})$ has a regular triple $(x,y,z)$. Write $K$ as
@@
where $|u|=4$, ${\rm Q}_{8}=\pm\{1,i,j,k\}$ with $i^{2}=j^{2}=k^{2}=ijk=-1$, and $\sigma,\tau\in{\sf Aut}({\rm Q}_{8})$ such that
$\sigma:i\rightarrow j\rightarrow k\,(\rightarrow i),\mbox{\,\ and\ \,}\tau:i\leftrightarrow-j,j\leftrightarrow-i,k\leftrightarrow-k.$
Then the subgroup $\langle\sigma\rangle{:}\langle\tau\rangle\cong{\rm S}_{3}$ contains three involutions, i.e., $\tau_{k}=\tau$,
$\tau_{i}:j\leftrightarrow-k,k\leftrightarrow-j,i\leftrightarrow-i,\mbox{\ and\ }\tau_{j}:i\leftrightarrow-k,k\leftrightarrow-i,j\leftrightarrow-j.$
By calculation, we conclude that $K$ contains the following involutions:
@@
$\lambda^{\lambda^{\prime}}=\lambda^{\tau_{\lambda}}=\lambda^{\lambda\tau_{\lambda}}=-\lambda$, and $\lambda^{\tau_{\lambda^{\prime}}}=-\lambda^{\prime\prime},\ \lambda^{\lambda^{\prime}\tau_{\lambda^{\prime}}}=-\lambda^{\tau_{\lambda^{\prime}}}=\lambda^{\prime\prime}$,
@@
$(\pm u\lambda)^{\pm\tau_{\lambda^{\prime}}}=\pm u(\lambda^{\pm\tau_{\lambda^{\prime}}})=\mp u\lambda^{\prime\prime},\ (\pm u\lambda)^{\pm\lambda^{\prime}\tau_{\lambda^{\prime}}}=\pm u(\lambda^{\pm\lambda^{\prime}\tau_{\lambda^{\prime}}})=\pm u\lambda^{\prime\prime}.$
@@
Let $H=\langle u,i,j,k\rangle\leqslant K$. Then $H\cong{\sf Z}_{4}\circ{\rm Q}_{8}$ is normal in $K$ and $K/H\cong{\rm S}_{3}$. Further, the involutions contained in $H$ are: $\{-1\}\cup\{\pm u\lambda\}$. Recall that $(x,y,z)$ is a regular triple for $G$. Thus, $\langle x,z\rangle\cong{\sf Z}_{2}^{2}$, and so $\langle x,z\rangle\cap H>1$, that is,
@@
By the claim given above, one of $x,z,xz$ has to be 1-1. Thus, $\overline{K}=K/\langle-1\rangle=\langle\overline{x},\overline{y}\rangle$ or $\langle\overline{y},\overline{z}\rangle$ is a dihedral group, which is impossible as $\overline{K}=K/\langle-1\rangle={\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\rm S}_{3})$. $\Box$
@@
Let $(\alpha,z)$ be a rotary pair for $G$. Since $G$ is not dihedral, $|\alpha|\neq 2$, so $\alpha\in\langle a,b\rangle$ and $\langle\alpha\rangle$ is normal in $G$. It follows that $\overline{G}=G/\langle\alpha\rangle=\langle\overline{z}\rangle={\sf Z}_{2}$, not possible. $\Box$
Now we are ready to prove Theorem 1.2.
@@
Let $\mathcal{M}$} be a map with square-free Euler characteristic, and let $G\leqslant{\sf Aut}{\mathcal{M}}$ be a solvable arc-transitive automorphism group of $\mathcal{M}$}. By Lemmas 2.1 and 2.3, the group $G$ satisfies Hypothesis 2.2, so that
G=(A:B):KG=(A{:}B){:}K
as described in Theorem 1.1.
@@
Recall that $\mathcal{M}$} as a $G$-arc-transitive map satisfies one of the following properties: (i) $G$-regular; (ii) $G$-vertex-reversing; (iii) $G$-vertex-rotary; respectively in each case, the group $G$ has: (i) a regular triple; (ii) a reversing triple; (iii) a rotary pair. In particular, $G$ is of even order, and so $K\neq 1$.
@@
First, suppose $K=G_{2}$. By Lemma 6.1, either $K$ has a regular triple/reversing triple/rotary pair as $G$ does, and then $K$ is one of the 2-groups given in Proposition 4.2; or $K$ is ${\sf Z}_{2}$, dihedral ${\rm D}_{2^{\ell}}$ or cyclic ${\sf Z}_{2^{\ell}}$ with $\ell\geqslant 2$, while ${\rm D}_{2^{\ell}}$ and ${\sf Z}_{2^{\ell}}$ already appear in Proposition 4.2.
@@
Then, suppose $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}G_{7}$ or $K=G_{\{2,3,7\}}={\sf Z}_{2}^{3}{:}(G_{7}{:}G_{3})$. In this case, $G$ has an odd factor group $G_{7}$ or $G_{7}{:}G_{3}$, respectively. By Lemma 6.1, G has neither regular triples nor reversing triples, so $G$ has a rotary pair, with each odd factor group cyclic. It follows that $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}{\sf Z}_{7^{\ell}}$.
@@
At last, suppose that $K=G_{\{2,3\}}$ is one of the groups given in 5.6 (2) - (3).
@@
(a) If $\mathcal{M}$} is $G$-vertex-reversing, then $G$ has a reversing triple, with each factor group of even order by Lemma 6.1. Hence, $K$ is one of the groups in Tables 1 - 2, excluding those corresponding to cases (1.3), (1.4), (1.7), (2.1) - (2.3) as they have an odd factor group ${\sf Z}_{3}$ or ${\sf Z}_{3^{\ell}}$.
@@
(b) If $\mathcal{M}$} is $G$-regular, then $G$ has a regular triple. In particular, $G$ has a reversing triple, and so $K$ is one of the groups in (a). Furthermore, by Lemma 6.2 the groups which have ${\rm Q}_{8}{:}{\rm S}_{3}$ or ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm S}_{3})$ as a factor group should be excluded. It follows that $K$ is one of: $X$, $X\times{\sf Z}_{2}$, ${\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ or $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$, where
$X\in\{{\rm S}_{4},\,{\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}},\,({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\,({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}\}.$
@@
(c) If $\mathcal{M}$} is $G$-vertex-rotary, then $G$ has a rotary pair, with each odd factor group cyclic by Lemma 6.1. So, either $K=X{:}K_{3}=X{:}{\sf Z}_{3^{\ell}}$ with $X\in\{{\sf Z}_{2}^{2},\,{\sf Z}_{2}^{3},\,{\rm Q}_{8},\,{\sf Z}_{4}\circ{\rm Q}_{8}\};$ or $K$ is one of the groups in Tables 1 - 2. For the former, $X\neq{\rm Q}_{8}$, as the only involution of ${\rm Q}_{8}{:}{\sf Z}_{3^{\ell}}$ lies in the center and so it does not have a rotary pair. For the latter, by Lemma 6.3 the groups corresponding to cases (1.5) and (1.6), which have $({\sf Z}_{3^{\ell}}\times{\sf Z}_{3}){:}{\sf Z}_{2}$ as a factor group, should be excluded.
In conclusion, $K$ is one of the groups stated in Theorem 1.2. $\Box$

### v3

@@ 4. $p$-Groups
Let $(x,y,z)$ be a regular triple for $G$. Then, $G=\langle x,y,z\rangle$ and $\langle x,z\rangle={\sf Z}_{2}^{2}$, so $\overline{G}=\langle\overline{x},\overline{y},\overline{z}\rangle$, $\overline{x}\,\overline{z}=\overline{z}\,\overline{x}$. Now, if $\overline{x},\overline{y},\overline{z}$ are involutions, and $\overline{x}\neq\overline{z}$, then $(\overline{x},\overline{y},\overline{z})$ is a regular triple for $\overline{G}$. Otherwise, either some of $\overline{x},\overline{y},\overline{z}$ equals the identity-element $\overline{1}$, or $\overline{x}=\overline{z}$, where, in both cases, $\overline{G}$ is generated by at most two involutions, so it is ${\sf Z}_{2}$ or dihedral. In particular, $\overline{G}$ is always of even order.
@@
Let $G={\rm GL}(2,3)$. Then neither $G$ nor ${\sf Z}_{4}\circ G$ has a regular triple.
@@
Suppose that $G\cong{\rm Q}_{8}{:}{\rm S}_{3}$ has a regular triple $(x,y,z)$. Then, $\langle x,z\rangle={\sf Z}_{2}^{2}$, so $\{x,z,xz\}\cap{\rm Q}_{8}\not$$\emptyset$, namely, one of $x,z,xz$ is the central involution 1-1 of $G$. It follows that $\overline{G}=G/\langle-1\rangle=\langle\overline{x},\overline{y}\rangle$ or $\langle\overline{y},\overline{z}\rangle$ is dihedral, and this contradicts $\overline{G}\cong{\sf Z}_{2}^{2}{:}{\rm S}_{3}={\rm S}_{4}$.
Suppose that $K={\sf Z}_{4}\circ G\cong{\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm S}_{3})$ has a regular triple $(x,y,z)$. Write $K$ as
@@
where $|u|=4$, ${\rm Q}_{8}=\pm\{1,i,j,k\}$ with $i^{2}=j^{2}=k^{2}=ijk=-1$, and $\sigma,\tau\in{\sf Aut}({\rm Q}_{8})$ act on ${\rm Q}_{8}$ as follows :
$\sigma:i\rightarrow j\rightarrow k\,(\rightarrow i)\mbox{\,\ and\,\ }\tau:i\leftrightarrow-j,j\leftrightarrow-i,k\leftrightarrow-k.$
Then, $\langle\sigma\rangle{:}\langle\tau\rangle={\rm S}_{3}$ contains three involutions, i.e., $\tau_{k}:=\tau$, and $\tau_{i}$, $\tau_{j}$ such that
$\tau_{i}:j\leftrightarrow-k,k\leftrightarrow-j,i\leftrightarrow-i\mbox{\,\ and\,\ }\tau_{j}:i\leftrightarrow-k,k\leftrightarrow-i,j\leftrightarrow-j.$
By direct calculation, we conclude that the involutions contained in $K$ are as follows :
@@
$\lambda^{\lambda^{\prime}}=\lambda^{\tau_{\lambda}}=\lambda^{\lambda\tau_{\lambda}}=-\lambda$ and $\lambda^{\tau_{\lambda^{\prime}}}=-\lambda^{\prime\prime},\ \lambda^{\lambda^{\prime}\tau_{\lambda^{\prime}}}=-\lambda^{\tau_{\lambda^{\prime}}}=\lambda^{\prime\prime}$,
@@
[…]
@@
Let $H=\langle u,i,j,k\rangle\leqslant K$. Then, $H\cong{\sf Z}_{4}\circ{\rm Q}_{8}$ is normal in $K$, $K/H\cong{\rm S}_{3}$, and the involutions contained in $H$ lie in $\{-1\}\cup\{\pm u\lambda\}$. Now, $(x,y,z)$ is a regular triple for $G$ such that $\langle x,z\rangle={\sf Z}_{2}^{2}$, so $\langle x,z\rangle\cap H>1$, that is,
@@
By the above claim, one of $x,z,xz$ has to be 1-1. It follows that $\overline{K}=K/\langle-1\rangle=\langle\overline{x},\overline{y}\rangle$ or $\langle\overline{y},\overline{z}\rangle$ is dihedral, and this contradicts $\overline{K}\cong{\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\rm S}_{3})={\sf Z}_{2}\times{\rm S}_{4}$. $\Box$
@@
Let $(\alpha,z)$ be a rotary pair for $G$. Then $G=\langle\alpha,z\rangle$ and $|z|=2$. Since $G$ is not dihedral, $|\alpha|\neq 2$, so $\alpha\in\langle a,b\rangle$ and $\langle\alpha\rangle$ is normal in $G$. It follows that $\overline{G}=G/\langle\alpha\rangle=\langle\overline{z}\rangle={\sf Z}_{2}$, not possible. $\Box$
Now, we are ready to prove Theorem 1.2.
@@
Let $\mathcal{M}$} be a map with square-free Euler characteristic, and let $G\leqslant{\sf Aut}{\mathcal{M}}$ be a solvable arc-transitive automorphism group of $\mathcal{M}$}. By Lemmas 2.1 and 2.3, the group $G$ satisfies Hypothesis 2.2, so by Theorem 1.1, we have :
G=H:K=(A:B):K,G=H{:}K=(A{:}B){:}K,
where $A$ is abelian, $B$ is nilpotent and $A\cap{\bf Z}(A{:}B)=1$. We thus only need to present $K$.
@@
Recall that $\mathcal{M}$} as a $G$-arc-transitive map satisfies one of the following properties : $G$-regular ; $G$-vertex-reversing ; $G$-vertex-rotary ; respectively, in each case, the group $G$ has : a regular triple ; a reversing triple ; a rotary pair. In particular, $G$ is of even order, so $K\neq 1$.
@@
(a) Suppose $K=G_{2}$. By Lemma 6.1, either $K$ has a regular triple/reversing triple/rotary pair as $G$ does, and then $K$ is one of the 2-groups given in Proposition 4.2; or $K$ is ${\sf Z}_{2}$, dihedral ${\rm D}_{2^{\ell}}$ or cyclic ${\sf Z}_{2^{\ell}}$ ($\ell\geqslant 2$), while ${\rm D}_{2^{\ell}}$ and ${\sf Z}_{2^{\ell}}$ already appear in Proposition 4.2.
This give the Theorem 1.2 (1).
(b) Suppose that $K$ is one of the groups given in Theorem 1.1 (ii) and (iv) :
${\sf Z}_{2}^{2}{:}G_{3},\ {\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}G_{3}),\ {\rm Q}_{8}{:}G_{3},\ {\sf Z}_{4}\circ_{{\sf Z}_{2}}({\rm Q}_{8}{:}G_{3}),\ {\sf Z}_{2}^{3}{:}G_{7},\ {\sf Z}_{2}^{3}{:}(G_{7}{:}G_{3}).$
@@
Then, $G$ has an odd factor group $G_{3}$, $G_{7}$ or $G_{7}{:}G_{3}$. By Lemma 6.1, $G$ has neither regular triples nor reversing triples, so $G$ has a rotary pair, with each odd factor group cyclic, namely, $G_{3}={\sf Z}_{3^{\ell}}$, $G_{7}={\sf Z}_{7^{\ell}}$ and $K\neq{\sf Z}_{2}^{3}{:}(G_{7}{:}G_{3})$. Thus, $K$ is one of :
${\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}}$,  ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})$,  ${\sf Z}_{4}\circ_{{\sf Z}_{2}}({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})$,  ${\sf Z}_{2}^{3}{:}{\sf Z}_{7^{\ell}},$
where ${\rm Q}_{8}{:}{\sf Z}_{3^{\ell}}$ is excluded, because it has the only involution lying in the center, and consequently, it does not have a rotary pair.
This gives the first half of Theorem 1.2 (4).
@@
(c) At last, suppose that $K$ is one of the groups given in Theorem 1.1 (iii) : $X$, ${\sf Z}_{2}\times X$, $Y$, ${\sf Z}_{4}\circ_{{\sf Z}_{2}}Y$ and $W$, where $X,Y,W$ are given in Table 1.
@@
If $\mathcal{M}$} is $G$-vertex-reversing, then $G$ has a reversing triple, with each factor group of even order by Lemma 6.1. Note that, the groups $X,Y$ located in rows $2,3,6$ of Table 1 have an odd factor group ${\sf Z}_{3}$ or ${\sf Z}_{3^{\ell}}$, so does the groups $W$ located in rows $1,2,5$. Thus, $X,Y,W$ can only taken from the others, namely,
$X\in\{{\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}},\,({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\,(({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})\times{\sf Z}_{3}){:}{\sf Z}_{2}\},$
$Y\in\{{\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}},\,({\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\sf Z}_{3})){:}{\sf Z}_{2},\,({\sf Z}_{3}\times({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}\}$, and
$W\in\{{\rm D}_{2.3^{\ell}}\times{\rm S}_{4},\,({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6},\,({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ({\rm Q}_{8}{:}{\rm S}_{3}),\,({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})\circ({\sf Z}_{3}{:}{\sf Z}_{4})\}$.
@@
If $\mathcal{M}$} is $G$-regular, then $G$ has a regular triple, so it has a reversing triple, and $K$ is one of the above groups. Furthermore, by Lemma 6.2, the groups which have ${\rm GL}_{2}(3)$ or ${\sf Z}_{4}\circ{\rm GL}_{2}(3)$ as a factor group should be excluded. Hence, $G\neq Y$, ${\sf Z}_{4}\circ_{{\sf Z}_{2}}Y$, $({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ({\rm Q}_{8}{:}{\rm S}_{3})$ or $({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})\circ({\sf Z}_{3}{:}{\sf Z}_{4})$.
This gives Theorem 1.2 (2) and (3).
@@
(c.3) If $\mathcal{M}$} is $G$-vertex-rotary, then $G$ has a rotary pair, so by Lemma 6.3, the following four groups, which have $({\sf Z}_{3^{\ell}}\times{\sf Z}_{3}){:}{\sf Z}_{2}$ as a factor group, should be excluded :
$({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2}$,  $({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$,  $({\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\sf Z}_{3})){:}{\sf Z}_{2}$,  $({\sf Z}_{3}\times({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$.
Finally, this completes Theorem 1.2 (4). $\Box$